OpenStudy (anonymous):

Help me please? :) A ball is dropped from the top of a 1,000 foot building. The height of the ball is half its original height after each bounce. What will the height of the ball be after 10 bounces?

OpenStudy (anonymous):

This is basically a half life equation, and the easiest way to go about it is dividing by two each time, so 1,000/2/2/2/2/2/2/2/2/2/2 and that's .9765625 feet I think.

OpenStudy (anonymous):

Thank you! So a decay in sorts? 1000 / 2^10

OpenStudy (anonymous):

yes h90) =1000 h(1) = 1000/2^1 = 500 h(2) = 1000/2^2 = 250 etc